Two cars A and B are moving in the same direction with velocities equal to 10 m/s and 5 m/s . B is ahead of car A and B has an acceleration of 2 m/s^2. The distane b/w A and B is 40 m . Find the minimum seperation b/w them.

Grade:12

1 Answers

san
10 years ago
velocity of b wrt a= velocity of b wrt ground - velocity of a wrt ground = 5 -10=-5m/s acc of b wrt a=acc of b wrt ground - acc of a wrt ground =2-0=2m/s now if we consider a person sitting in a he will see first coming towards him till its velocity becomes zero [ as v b wrt a and acc b wrt a are in opposite dir] nd then b moves forward away from a so using v2 - u2=2as 0-25=-4s 6.25=s this is dis travelled by b acc to a till its velocity becomes zero[nd then it moves forward] as we r seeing b frm a we say a is at rest so min dis between the = 40 -6.25=33.75m